### v1

# GS-ProCams: Gaussian Splatting-based Projector-Camera Systems

Abstract
We present GS-ProCams, the first Gaussian Splatting-based framework for projector-camera systems (ProCams). GS-ProCams significantly enhances the efficiency of projection mapping (PM) that requires establishing geometric and radiometric mappings between the projector and the camera. Previous CNN-based ProCams are constrained to a specific viewpoint, limiting their applicability to novel perspectives. In contrast, NeRF-based ProCams support view-agnostic projection mapping, however, they require an additional colocated light source and demand significant computational and memory resources. To address this issue, we propose GS-ProCams that employs 2D Gaussian for scene representations, and enables efficient view-agnostic ProCams applications. In particular, we explicitly model the complex geometric and photometric mappings of ProCams using projector responses, the target surface’s geometry and materials represented by Gaussians, and global illumination component. Then, we employ differentiable physically-based rendering to jointly estimate them from captured multi-view projections. Compared to state-of-the-art NeRF-based methods, our GS-ProCams eliminates the need for additional devices, achieving superior ProCams simulation quality. It is also 600 times faster and uses only 1/10 of the GPU memory.
Figure 1: GS-ProCams. We represent the scene using 2D Gaussian, allowing us to efficiently model the geometric and photometric mappings of ProCams. By projecting various patterns onto the target surface and capturing images from multiple viewpoints using a camera, we achieve high-fidelity ProCams simulation while estimating the target surface’s geometry and materials. After training on a scene, GS-ProCams can achieve efficient view-agnostic (a) projector compensation, (b) projection-based diminished reality [29, 45] (with inpainting models [65]), and (c) projection mapping [28] (with diffusion models [53, 21]).
1 Introduction
Projector-camera systems (ProCams) are integrated systems that combine a projector and a camera, and have emerged as indispensable tools in computer vision [60, 43, 5, 76, 3, 44, 61], spatial augmented reality (SAR)/projection mapping (PM) [66, 57, 48, 30, 13, 56, 62, 67, 22, 12, 14, 15], providing exceptional experiences in immersive interaction, artistic creation and exhibition, as well as industrial design and manufacturing, among other domains.
Geometric and photometric mappings between the projector and the camera, which are foundational to ProCams, serve as cornerstones for numerous applications. For example, structured light utilizes the geometric and photometric constraints of ProCams to estimate the object geometry and materials [55, 54, 18, 39]. Furthermore, by ProCams geometric correction [6, 69, 52, 78, 75] and photometric compensation [74, 37, 19, 2, 23, 47], full projector compensation can be achieved to rectify projections on non-planar and textured surfaces, ensuring optimal visual effects.
Although recent advancements in Convolutional Neural Network-based schemes [33, 24, 32, 26, 72, 73] have greatly improved the convenience and quality of projection mapping, these methods are usually limited to a specific viewpoint, making them incapable of direct application to novel viewpoints. Neural fields-based method [14] utilize multilayer perceptrons (MLPs) [7, 64, 83] for projection mapping. Despite their capability for novel viewpoint ProCams applications, this method requires an additional co-located light and a dark room (i.e., assumes no ambient light). Moreover, it suffers from low computational and memory efficiency.
In this paper, we propose GS-ProCams for efficient view-agnostic projection mapping. In particular, we first utilize 2D Gaussians [27] as ProCams scene representation, since the original 2D Gaussians do not explicitly capture Bidirectional Reflectance Distribution Function (BRDF), we augment each Gaussian point in space with BRDF attributes. Then, we capture images of the target surface within the scene from multiple camera viewpoints while using a fixed projector to project various patterns onto it. Afterward, we employ structure-from-motion (SfM) [63, 58] to determine the poses and intrinsic parameters of the projector and the cameras. Subsequently, we jointly reconstruct the target surface and the projector’s response by differentiable physically-based rendering across each viewpoint. Finally, we apply this trained GS-ProCams model to various view-agnostic ProCams applications, such as projection mapping and projector compensation, as shown in Fig. 1. Comprehensive experiments on ProCams simulation and projector compensation show clear advantages of our GS-ProCams over previous methods.
In summary, our primary contributions include:
  • We present GS-ProCams, the first Gaussian Splatting-based framework for projector-camera systems. It integrates 2D Gaussians and physically-based rendering to efficiently model the geometric and photometric mappings of ProCams.
  • Compared to NeRF-based ProCams, GS-ProCams eliminates the need for additional devices, can operate under ambient room light, and is 600 times faster while using only 1/10 of the GPU memory.
  • GS-ProCams can be applied to multiple view-agnostic ProCams applications simultaneously, including projection mapping and projector compensation.
  • We introduce a view-agnostic ProCams benchmark encompassing various viewpoints, textured surfaces, and ambient lighting conditions for training and evaluating ProCams models. This benchmark is designed to foster future research in the field.
2 Related Work
2.1 ProCams simulation
ProCams simulation aims to model the real-world photometric and geometric mappings of projector-camera systems (ProCams) within a computational space. Once these mappings are established, the simulation synthesizes camera images corresponding to various projections. ProCams simulation facilitates image-based relighting [74, 24, 14], shape reconstruction [60, 43, 24], and projector compensation [32, 26]. Early methods describe the pixel mapping from projectors to cameras using the Light Transport Matrix (LTM) [59, 49, 50, 74]. However, the reconstruction quality of LTM, which is based on the linear relationship between pixels, generally depends on dense sampling and specialized computational designs of large matrices [10, 11, 59, 74, 49, 70]. By employing neural rendering, DeProCams [24] generates photorealistic simulations of ProCams with fewer image samples by incorporating inherent epipolar geometric and simplified photometric constraints under a fixed viewpoint. However, a notable limitation of this approach is its reliance on manual calibration of ProCams using a checkerboard pattern [25], rather than leveraging data priors [58]. Although Park et al. [51] examine the geometric and photometric mappings in ProCams using a differentiable rendering scheme, it remains limited to specific view information and relies on an RGB-D camera. Additionally, the method depends heavily on pixel-level bias corrections. Erel et al. [14] integrate projectors into Neural Reflectance Fields [7, 64, 83] to enable view-agnostic projection mapping within 3D scenes. Despite this progress [14], the influence of other light sources within the scene is neglected, and a light source co-located with the camera is required, thereby causing limited applicability. We summarize the characteristics of technologies for ProCams simulation to provide an intuitive comparison with our approach in Tab. 1.
Table 1: Comparison of representative schemes for ProCams and proposed GS-ProCams. GS-ProCams is a view-agnostic method, which enables multi-view applications. It also simultaneously addresses geometric and photometric mappings under ambient light, using only an RGB camera and a projector.
Method View-agnostic W/o manual W/o extra Ambient
method calibration devices light
LTM [74]
CNN-based [24]
Depth-prior DR [51]
NeRF-like DR [14]
GS-ProCams (Ours)
CNN: Convolutional Neural Network; DR: Differentiable Rendering
2.2 Projector compensation
Projector compensation aims to adjust projections for non-planar and textured surfaces to achieve the best visual display effect by optimizing the projector’s input patterns. Typically, projector compensation can be conducted separately through geometric correction [6, 69, 52, 78, 75] and radiometric compensation [19, 37, 2, 23, 47]. Upon completing the ProCams simulation, the compensated projection patterns can be obtained by multiplying the desired images with the inverse LTM [74]. Differentiable ProCams [24, 51, 14] generate projection patterns by making the synthesized scene closely approximate the desired images. On the other hand, comprehensive CNN-based architectures enable end-to-end projector compensation by ingeniously integrating geometric correction estimation represented by the warping grid [26, 72] or optical flow [73, 33, 40] with another component for photometric compensation. However, these methods are typically constrained to a specific viewpoint and smooth, uniform surfaces. Erel et al. [14] achieve projector compensation from novel viewpoints by integrating a projector into Neural Reflectance Fields [7, 64, 83]. Despite that, this method, which relies on ray-casting and MLPs, poses challenges to computational and memory resources. Additionally, this method assumes no ambient light in the scene.
2.3 Gaussian splatting and inverse rendering
3D Gaussian Splatting (3DGS) [35] has quickly sparked widespread interest due to its real-time differentiable rendering capability. While achieving rendering quality comparable to the state-of-the-art methods in novel view synthesis [42, 4, 77, 46, 16], it significantly enhances the efficiency of training and rendering. Furthermore, inverse rendering schemes extend 3DGS for materials and ambient lighting estimations [31, 17, 41].
While 3DGS [35] focuses primarily on efficient image rendering capabilities, reconstructing precise geometric shapes directly from Gaussian points poses significant challenges [27, 9, 80, 20, 81, 68]. In contrast, 2D Gaussian Splatting (2DGS) [27] uses 2D Gaussian primitives to ensure geometric consistency across multiple viewpoints. We enhance 2DGS with BRDF parameterization to capture the interactions between projector lighting and surfaces. We first use 2DGS to generate depth and materials maps in the camera frame and then perform shading calculations. The entire process is fully explicit and differentiable.
Figure 2: GS-ProCams setup. The intersection point $\mathbf{x}_{s}$ of the ray from the camera and the target surface is determined by 2D Gaussians. The direct light from the projector illuminates this point.
Figure 3: Pipeline. GS-ProCams uses 2D Gaussians to represent the scene, then obtains the geometry and materials of the target surface, as well as the global illumination component, through differentiable splatting techniques [27]. By using various projection patterns and capturing images from different viewpoints, GS-ProCams jointly optimizes the explicit projector responses and attributes of the 2D Gaussian points listed on the left through physically-based differentiable rendering.
3 GS-ProCams
3.1 Preliminary
2DGS [27] employs explicit 2D Gaussians to represent scene geometry. The following primary components distinguish each 2D Gaussian point: its central point $\mathbf{p}_{k}$, two principal tangential vectors $\mathbf{t}_{u}$ and $\mathbf{t}_{v}$, and a scaling vector $\mathbf{s}=(s_{u},s_{v})$ which regulates the variances of the 2D Gaussian distribution. Consequently, 2D Gaussians inherently include inherent normals defined by the above tangential vectors as $\mathbf{t}_{w}=\mathbf{t}_{u}\times\mathbf{t}_{v}$. In world space, every 2D Gaussian is defined in a local tangent plane and parameterized as:
${P(u,v)=\mathbf{p}_{k}+s_{u}\mathbf{t}_{u}u+s_{v}\mathbf{t}_{v}v.}$ (1)
Then, the 2D Gaussian value for a point $\mathbf{u}=(u,v)$ in its local space can be evaluated by the standard Gaussian:
$G(\mathbf{u})=\text{exp}\Big{(}-\frac{u^{2}+v^{2}}{2}\Big{)}.$ (2)
Additionally, each 2D Gaussian primitive possesses an opacity $o$ and a view-dependent appearance $\mathbf{c}$ represented using spherical harmonics (SH). Finally, given the viewpoint-based color $\mathbf{c}_{i}$ of each Gaussian point, the pixel color $\mathbf{C}$ is obtained through alpha blending with $K$ tile-based Gaussians arranged according to the depth values of their center locations:
$\mathbf{C}(\mathbf{x})=\sum_{i\in K}W_{i}c_{i},\;\;W_{i}=\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (3)
where $\alpha=o\cdot\hat{G}(u(\mathbf{x}))$. Here, $\hat{G}(u(\mathbf{x}))$ is the result of applying a low-pass filter to ${G}(u(\mathbf{x}))$, and $\mathbf{x}$ denotes a homogeneous ray that originates from the camera and traverses the UV space.
The orientation of a 2D Gaussian can be construed as a rotation matrix $\mathbf{R}=[\mathbf{t}_{u},\mathbf{t}_{v},\mathbf{t}_{w}]$ and be represented in the form of a unit quaternion $\mathbf{q}$ during optimization. In summary, each 2D Gaussian is parameterized by several learnable parameters as $\{\mathbf{p},\mathbf{q},\mathbf{s},o,\mathbf{c}\}$. Furthermore, the designed density control mechanism enables adaptive regulation of the total number of points and the density within a unit volume.
3.2 ProCams modeling
The core of our proposed GS-ProCams is establishing the geometric and photometric mappings of projector-camera systems (ProCams) using Gaussian splatting, thus achieving high-fidelity ProCams simulation. Consider a point $\mathbf{x}_{s}$ on the target surface within a scene, which is captured by the camera as $\mathbf{x}_{c}$ in the camera frame, as illustrated in Fig. 2. In the absence of occlusion, this point is illuminated by $\mathbf{x}_{p}$ in the projector frame. Furthermore, $\mathbf{x}_{s}$ is affected by ambient and complex indirect lighting, i.e., global illumination. Taking these factors into consideration, We utilize the classical rendering equation [34] to describe the relationship between outgoing and incoming light at point $\mathbf{x}_{s}$ as follows:
[…]
where $\Omega$ represents the upper hemisphere centered at $\mathbf{x_{s}}$, with the normal vector $\mathbf{n}_{s}$ and $f$ denotes the BRDF. We employ 2D Gaussians described in 2D Gaussian Splatting (2DGS) [27] as foundational primitives, extending their original attributes by BRDF parameterization to capture the interactions between the light and the target surface. Building upon this, We capture a series of images of the surface $\{\mathbf{I}_{c}^{m}\}_{m=1}^{M}$ from multiple viewpoints, which is illuminated by known projector input patterns $\{\mathbf{I}_{p}^{n}\}_{n=1}^{N}$. Then, we simultaneously estimate the surface’s unknown geometric and material properties utilizing physically-based rendering (PBR). We illustrate the pipeline of our proposed GS-ProCams in Fig. 3.
3.3 Attributes from Gaussians
As shown in Fig. 3, we first derive attributes of the target surface from 2D Gaussians. The attributes can generally be divided into three categories: (1) the depth and inherent normals from geometry attributes of Gaussians; (2) the BRDF attributes for material reflectance; and (3) the global illumination component, approximated by RGB residuals. Notably, we utilize normals derived from depth gradients in the shading computations, while use inherent normals to regularize geometric consistency.
Depth. The intersections between camera rays and the surface are determined by their depths in the camera frame. We use the same approach as 2DGS [27] and render the mean depth $D$ by Eq. 3 to represent the ray-surface intersection as follows:
$D=\sum_{i\in K}\frac{W_{i}z_{i}}{\sum_{i\in K}W_{i}},$ (5)
where $\mathbf{x}$ is the homogeneous ray corresponding to the camera pixel $\mathbf{x}_{c}=(x,y)$, and $z_{i}$ denotes the depth of the intersection between $\mathbf{x}$ and the $i$-th Gaussian planes.
Normal. To enhance depth consistency and surface smoothness, we utilize normals derived from viewpoint depth rather than the inherent normals aligned with the directions of their highest density gradients as the normal map $\mathbf{N}$:
$\mathbf{N}(x,y)=\frac{\nabla_{x}\mathbf{d}\times\nabla_{y}\mathbf{d}}{\left|\nabla_{x}\mathbf{d}\times\nabla_{y}\mathbf{d}\right|},$ (6)
where x and y denote the coordinates of the pixel $\mathbf{x}_{c}$ in the depth map $D$.
BRDF attributes. To enable the light-surface interactions, we augment each original 2D Gaussian point [27] with additional learnable attributes: albedo $\mathbf{b}\in[0,1]^{3}$ and roughness $r\in[0,1]$. Then we splat these attributes onto the camera frame by utilizing the efficient differentiable rasterization method described by Eq. 3:
$\{\mathbf{B},R\}=\sum_{i\in K}W_{i}\{\mathbf{b}_{i},r_{i}\},$ (7)
where $\mathbf{B}$ and $R$ represent the maps of albedo and roughness after rasterization in a specified viewpoint.
3.4 Lighting
In ProCams settings, light usually comes from the environment and the projector. The projector pixels act like high-power spotlight light sources that dominate the light contribution. Therefore, we model the incident light at $\mathbf{x}_{s}$ as the sum of projector direct light, denoted as $L_{p}$, and the global illumination, denoted as $L_{g}$.
Projector direct light. Denote the projector gamma as $\gamma_{p}$, which converts standard sRGB color into linear RGB space. Denote the projector’s luminous power as $G_{p}$, and assume constant light attenuation. Additionally, due to the projector’s optical characteristics, e.g., point spread function (PSF) [38], the projected light suffers from defocus. Therefore, we use a learnable $5\times 5$ kernel $\kappa$ to model the projector’s PSF. Consequently, the projector emitted direct light $L_{p}$ can be expressed as:
$L_{p}(\mathbf{x}_{s})=\kappa\ast\big{(}G_{p}\cdot\mathbf{I}_{p}^{\gamma_{p}}(\mathbf{x}_{p})\big{)},$ (8)
where $\mathbf{x}_{p}$ is the projector pixel that directly illuminates the surface point $\mathbf{x}_{s}$, as shown in Fig. 2, and $\ast$ is the convolution operator. Denote the projector projection matrix as $\mathbf{M}_{p}$, the relationship between $\mathbf{x}_{p}$ and $\mathbf{x}_{s}$ is given by
$\overline{\mathbf{x}}_{p}=\mathbf{M}_{p}\overline{\mathbf{x}}_{s},$ (9)
where $\overline{\mathbf{x}}_{p}$ and $\overline{\mathbf{x}}_{s}$ denote the homogeneous coordinates of $\mathbf{x}_{p}$ and $\mathbf{x}_{s}$, respectively. Then, according to Eq. 4, the outgoing/reflected direct light at $\mathbf{x}_{s}$ can be computed by:
$\mathbf{C}_{p}(\omega_{o},\mathbf{x}_{s})=f(\omega_{o},\omega_{p},\mathbf{x}_{s})L_{p}(\omega_{p},\mathbf{x}_{s})(\omega_{p}\cdot\mathbf{n}_{s}),$ (10)
where the normal $\mathbf{n}_{s}$ of the surface point $\mathbf{x}_{s}$ is determined by Eq. 6. We approximate the BRDF $f$ in Eq. 10 using a simplified Disney BRDF model [8].
Global illumination. In most ProCams settings, projector direct light dominates the illumination, therefore, we utilize a simplified generalized estimate to replace the precise yet complex calculations for efficiency. Inspired by 3DGS [35] and 2DGS [27], we approximate the global illumination radiance as a view-dependent color $\mathbf{c}_{r}(\omega_{o})$ by maintaining a set of coefficients of the spherical harmonics for each Gaussian point and then rasterize the color using alpha blending according to Eq. 3 as follows:
$\mathbf{C}_{g}=\sum_{i\in K}W_{i}\mathbf{c}_{g_{i}}.$ (11)
Finally, the camera-captured color is given by
$\hat{\mathbf{I}}_{c}=\text{clamp}\left((\mathbf{C}_{p}+\mathbf{C}_{g})^{\frac{1}{\gamma_{c}}},0,1\right),$ (12)
where $\gamma_{c}$ is a gamma tone mapping function [1].
Figure 4: Visual comparisons of ProCams simulation on the Nepmap synthetic dataset [14]. The first column displays an object from a novel viewpoint, the second column shows the object under a novel projection pattern, and the third and fourth columns present the results of two methods. We present two of the four scenes here as examples. Compared to Nepmap, our model exhibits finer details and more realistic color. Moreover, our method outperforms Nepmap on computation and memory efficiency by a significant margin (Table 2). See supplementary material for more results.
3.5 Optimization
As shown in Fig. 3, the optimizable parameters associated with each Gaussian point are $\{\mathbf{p},\mathbf{q},\mathbf{s},o,\mathbf{c}_{g},\mathbf{b},r\}$. The optimizable parameters of projector responses are $\{\gamma_{p},G_{p},\kappa\}$. We impose the following constraints to optimize the GS-ProCams parameters.
Photometric loss. In line with 3DGS [35], we consider the combined loss function of $\mathcal{L}_{1}$ and D-SSIM $\mathcal{L}_{D-SSIM}$ to minimize the photometric difference between the simulated result $\hat{\mathbf{I}}_{c}$ and camera captured ground truth $\mathbf{I}_{c}$ and set the weight $\lambda=0.2$:
$\mathcal{L}_{c}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{D-SSIM}.$ (13)
Geometric regularization. We also employ depth distortion loss and normal consistency loss [27] to encourage the 2D primitives to align with the thin surfaces closely:
$\mathcal{L}_{d}=\sum_{i,j}w_{i}w_{j}\left|z_{i}-z_{j}\right|,\quad\mathcal{L}_{n}=\sum_{i}w_{i}(1-\mathbf{n}_{i}^{\top}\mathbf{N}),$ (14)
where $\mathbf{n}_{i}$ is the inherent normal of the the intersection between $\mathbf{x}$ and the $i$-th splat.
Materials regularization. In this paper, we estimate surface materials from realistic scenes where roughness maps are anticipated to exhibit smooth characteristics in regions with uniform color [79]. Consequently, we define a bilateral smoothness term, guided by the albedo, as follows:
$\mathcal{L}_{material}={\left\|\nabla{R}\right\|}\text{exp}(-{\left\|\nabla{\mathbf{B}}\right\|}).$ (15)
Note that we detach the gradient propagation of the albedo $\mathbf{B}$ before computing this loss.
Mask entropy. Most projector applications focus only on regions of interest, such as the projector’s FOV [26] or objects [14], typically using binary masks to indicate these areas. In light of this, We introduce a cross-entropy constraint [71] to regularize the optimization for obtaining opaque surfaces by forcing the accumulated opacity $O=\sum_{i\in K}W_{i}$ to be aligned with the mask $O_{m}$. The constraint is defined as:
$\mathcal{L}_{entropy}=-O_{m}\log O-(1-O_{m})\log(1-O),$ (16)
Total Loss. Finally, we train GS-ProCams with the following loss function:
$\mathcal{L}=\mathcal{L}_{c}+\lambda_{d}\mathcal{L}_{d}+\lambda_{n}\mathcal{L}_{n}+\lambda_{m}\mathcal{L}_{material}+\lambda_{e}\mathcal{L}_{entropy},$ (17)
where $\lambda_{d}$, $\lambda_{n}$, $\lambda_{m}$, and $\lambda_{e}$ represent the weights for the respective loss terms, which are set to $1000$, $0.05$, $0.002$ and $0.1$, respectively. We jointly optimize the GS-ProCams parameters using Adam optimizer [36] for 30,000 training steps.
4 Experiments
To demonstrate the advantages of our approach, we compared it with several state-of-the-art ProCams simulation and/or projector compensation methods [14, 24, 26] in Sec. 4.1 and Sec. 4.2. In our experiments, we employed Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [82] to evaluate the ProCams simulation, and projector compensation quality. Additionally, we statistically examined the memory requirements, training time, and inference frames per second (FPS) on an NVIDIA A800 GPU.
4.1 Synthetic dataset
We compared our GS-ProCams with the state-of-the-art NeRF-based ProCams simulation approach [14], using their synthetic dataset to evaluate the ProCams simulation quality. We refer to this method as Nepmap for convenience. We utilized the four synthetic scenes released by Nepmap and their pre-trained models to compare 36 novel viewpoints and novel projection patterns for each scene. Quantitative results are presented in Tab. 2, and qualitative comparisons are illustrated in Fig. 4. The results indicate that our model not only outperforms Nepmap in terms of quality but also shows a significant improvement in efficiency, i.e., about 600 times faster and only uses 1/10 of GPU memory. Furthermore, our model does not require additional devices and takes ambient light into account as listed in Tab. 1.
Table 2: ProCams simulation results on the Nepmap synthetic dataset [14]. Results are averaged over 4 different scenes.
[…]
Nepmap [14] 27.23 0.945 0.061 $>3\text{hrs}$ 0.36 34.88 GB
GS-ProCams 33.07 0.974 0.032 9.3 min 222 02.02 GB
Figure 5: Visual comparisons of ProCams simulation on real-world benchmark dataset. We compare with DeProCams [24] (view-specific) at the last training viewpoint (first row) and at the novel viewpoint that our GS-ProCams has not previously seen (second row). Note that DeProCams (view-specific) necessitates repeated data acquisition and training for the novel viewpoint, whereas our approach can generalize to novel viewpoints. The first column displays the projector input pattern, the second column shows the surface (no superimposed projection) from the viewpoint, and the subsequent columns sequentially illustrate the camera-captured ground truth and the simulations results. See supplementary material for more results.
Figure 6: Intermediate results. GS-ProCams uses interpretable parameters to establish the geometric and photometric mappings of ProCams, jointly estimating them while achieving view-agnostic ProCams simulation. Here, “$L_{p}$” is in the linear color space. “Albedo” and “Global illumination” are converted from linear to sRGB color space for better visual experience.
Table 3: ProCams simulation comparisons on real-world benchmark dataset. Results are averaged over 10 different scenes. Note that DeProCams needs to additionally capture training data and retraining for novel viewpoints, while our GS-ProCams waives this requirement.
Method Trained Viewpoint Unseen Novel Viewpoint
[…]
DeProCams11footnotemark: 1 [24] 28.66 0.907 0.109 04.61 min 27.63 0.886 0.134 4.62 min 26.73 GB
w/o PSF (Ours)22footnotemark: 2 25.76 0.887 0.104 15.86 min 24.83 0.837 0.121 0 03.38 GB
GS-ProCams (Ours)22footnotemark: 2 27.22 0.911 0.083 15.69 min 26.11 0.859 0.100 0 03.49 GB
11footnotemark: 1 View-specific method; 22footnotemark: 2 View-agnostic method
4.2 Real-world benchmark dataset
We collected a benchmark dataset in real indoor lighting environments. Our projector-camera system consists of a Panasonic Lumix DC-S5 camera and an EPSON CB-965 projector, and an RGB fill light was employed to create various ambient lighting conditions. It is worth noting that unlike Nepmap [14], our method does not require additional light sources to work, the RGB fill light is only used to change ambient lighting conditions. The projector and camera resolutions are both set to $800\times 800$.
We projected multiple patterns for each viewpoint, including a fixed pattern to enhance surface texture for the SfM process [58] and a natural image for model training. We followed the methodology outlined in [26] to obtain projector FOV masks. In our experimental configuration, we projected natural patterns across 48 viewpoints. At the 48th viewpoint, an additional 48 natural patterns were projected for validation. Subsequently, we randomly repositioned the camera and projected 48 patterns once more to evaluate the applicability of our method from this novel viewpoint. See supplementary material for more details.
Because Nepmap [14] requires a light source to be co-located with the camera and only works for a dark room, we cannot compare with it on our ProCams settings. Instead, we compare with two state-of-the-art view-specific methods [24, 26].
Figure 7: Visual comparisons of projector compensation on real-world benchmark dataset. Similar to ProCams simulation (Fig. 5), the camera-captured results from different viewpoints within the scene are presented in two rows. Our model is applied directly to novel viewpoints without retraining in the second row, while the other two view-specific methods necessitate repeated data capturing and training. The first three columns respectively illustrate the surface appearance from the specified viewpoint, the uncompensated appearance, and the expected appearance. The last three columns display the compensated results from different models. Note that our method effectively compensates for geometry and photometry when applied to the novel viewpoint. See supplementary material for more results.
Table 4: Projector compensation results on real-world dataset. Results are averaged over 10 different scenes. Our approach is directly applied to the unseen novel viewpoint without retraining, while the other two methods must be retrained.
Method Trained Viewpoint Unseen Novel Viewpoint
[…]
DeProCams11footnotemark: 1 [24] 22.78 0.816 0.189 04.61 min 22.35 0.796 0.217 04.62 min 26.73 GB
CompenNeSt++11footnotemark: 1 [26] 24.02 0.834 0.161 13.25 min 24.05 0.830 0.173 13.21 min 30.79 GB
GS-ProCams (Ours)22footnotemark: 2 23.91 0.828 0.173 15.69 min 23.66 0.807 0.194 0 03.49 GB
Uncompensated 14.80 0.669 0.281 - 14.57 0.641 0.303 - -
11footnotemark: 1 View-specific method; 22footnotemark: 2 View-agnostic method
Figure 8: Ablation study of PSF estimation. We compare GS-ProCams with a degraded version that ignores projector PSF. Note that the result of “w/o PSF” shown in the zoomed-in patches exhibits unrealistically sharp projections.
ProCams simulation. We use the state-of-the-art method, DeProCams [24] (view-specific), as a baseline to evaluate the effectiveness of our approach (view-agnostic) at the unseen novel viewpoint. We present quantitative comparisons in Tab. 3 and illustrate qualitative findings with a scene example in Fig. 5. For the trained viewpoint, our GS-ProCams outperforms DeProCams, due to our GS-based formulation. For the unseen novel viewpoint, DeProCams shows better simulation quality, because it is view-specific and is trained on this viewpoint, while our GS-ProCams has never seen this viewpoint. Our method highlights its advantages compared to the view-specific approach. First, our method facilitates the synthesis of surface appearances with novel projections under novel viewpoints. Second, our method demonstrates a significant efficiency advantage in memory usage. Additionally, as previously noted, GS-ProCams effectively solves the view-agnostic projection mapping problem by physically-based rendering, while deriving interpretable quantities from multi-view images as shown in Fig. 6.
Projector compensation. Our proposed method facilitates ProCams simulation through a fully differentiable process, enabling the optimization of projection patterns to align rendered images with the desired effects. Once trained, our method allows for direct projector compensation from novel viewpoints without retraining or fine-tuning. We compare with two state-of-the-art projector compensation methods CompenNeSt++ [26] (view-specific) and DeProCams [24] (view-specific). The quantitative results are summarized in Tab. 4, while the visual comparisons are illustrated in Fig. 7. The results show that our method outperforms DeProCams under the trained viewpoint and is comparable to CompenNeSt++. For the novel viewpoint, there is a gap between our model and CompenNeSt++, but it still outperforms DeProCams. This is because both CompenNeSt++ and DeProCams are retrained for the novel viewpoint, but our GS-ProCams is directly applied to the novel viewpoint without retraining. The view-agnostic projector compensation and low memory requirements demonstrate our significant advantages in practical applications.
4.3 Ablation study
We conducted ablation study on the real-world benchmark dataset to validate the effectiveness of the projector’s PSF estimation (Eq. 8). We denote GS-ProCams without PSF as w/o PSF. The quantitative and qualitative results are shown in Tab. 3 and Fig. 8, respectively. Clearly, without estimating the projector’s PSF, the synthesized results ignores projector defocus, and exhibit unrealistically sharp projections, while GS-ProCams results are much closer to the camera-captured ground truth.
5 Conclusion
We introduce GS-ProCams, the first Gaussian Splatting-based framework for projector-camera systems (ProCams). This framework enables efficient novel viewpoint ProCams simulation and projector compensation without retraining. Compared with NeRF-based ProCams, our GS-ProCams waives the need for additional devices, and achieves higher ProCams simulation quality. Moreover, it is 600 times faster and uses only 1/10 of the GPU memory.
Limitations and future work. GS-ProCams has limitations in regions with strong specular reflections. More intricate photometric mapping calculations and optimization strategies are expected to further improve the framework’s performance, how to balance the simulation quality and efficiency is worth exploring in future work.
Supplementary Material
In the supplementary material, we provide a detailed description of the real-world data acquisition process in Appendix A, further implementation details of GS-ProCams in Appendix B, and additional qualitative results from experiments on ProCams simulation and projector compensation in Appendix C.
Appendix A Real-world data acquisition
We use an RGB camera to acquire data in indoor lighting environments. Since a fixed projector typically illuminates a specific region of the target, we randomly position multiple camera viewpoints along the hemispherical surface surrounding the target and capture data sequentially.
We employ the pinhole camera model to represent both the projector and the camera viewpoints. A digital speckle pattern as illustrated in Fig. 9 is projected for each viewpoint, enabling us to obtain the locations and rotations of the projector and the camera viewpoints using COLMAP [58]. When applying the trained GS-ProCams to a previously unseen viewpoint, we only need to capture an image from this viewpoint with the same projection and then register it into the existing COLMAP model.
Appendix B Implementation details
B.1 BRDF modeling
We use a simplified Disney BRDF model [8] to approximate the BRDF $f$. Specifically, the BRDF $f$ is composed of a diffuse term $f_{d}=\frac{\mathbf{B}}{\pi}$ and a specular term:
$f_{s}(\omega_{o},\omega_{p})=\frac{DFG}{4(\mathbf{N}\cdot\omega_{p})\cdot(\mathbf{N}\cdot\omega_{o})},$ (18)
where $D$, $F$, and $G$ denote the microfacet distribution function, Fresnel reflection and geometry factor. Define the half vector between $\omega_{o}$ and $\omega_{p}$ as $\mathbf{h}$, the normal distribution function $D$ is evaluated as follows:
$D=\frac{R^{4}}{\pi((\mathbf{N}\cdot\mathbf{h})^{2}(R^{4}-1)+1)^{2}},$ (19)
and the Fresnel reflection term is approximated as:
$F=0.04+(1-0.04)2^{(-5.55473(\omega_{o}\cdot\mathbf{h})-6.98316)(\omega_{o}\cdot\mathbf{h})}.$ (20)
Define $k=\frac{(R+1)^{2}}{8}$, the geometry factor is given by:
$G=\frac{(\mathbf{N}\cdot\omega_{p})(\mathbf{N}\cdot\omega_{o})}{(\mathbf{N}\cdot\omega_{p}\cdot(1-k)+k)(\mathbf{N}\cdot\omega_{o}\cdot(1-k)+k)}.$ (21)
Figure 9: Images registration. We projected a fixed pattern (left) and captured an image using the camera (right) for each viewpoint, thereby determining the positions of the projector and camera viewpoints, using COLMAP [58].
B.2 Applications
As the principal manuscript outlines, our approach constitutes an efficient and fully differentiable framework for ProCams applications. In this section, we provide further details regarding the implementation.
Projector compensation. Our method facilitates view-agnostic projector compensation, as demonstrated in our experiments. Once the ProCams simulation is completed, GS-ProCams represents projector compensation as an inverse rendering problem. Given the desired appearance in the camera frame as $\mathbf{I}_{d}$, our framework allows for optimizing a virtual projector input pattern $\hat{\mathbf{I}}_{p}$ to make the simulation result $\hat{\mathbf{I}}_{c}$ in the virtual space align with $\mathbf{I}_{d}$ as following:
$\hat{\mathbf{I}}^{*}_{p}=\arg\min_{\hat{\mathbf{I}}_{p}}\mathcal{L}(\hat{\mathbf{I}}_{c},\mathbf{I}_{d}),$ (22)
where $\hat{\mathbf{I}}^{*}_{p}$ is the compensated projector input for the desired visual display effect.
More applications. We present some additional applications of GS-ProCams as shown in Fig. 1, such as projection-based diminished reality and text-driven projection mapping, which are achieved by integrating GS-ProCams with popular large generative models. Specifically, for projection-based diminished reality, we eliminate objects captured in the camera frame using an inpainting model [65] and then apply the techniques as described in Sec. B.2 to obtain the projector input pattern, which can act on reality to remove the target objects visually. Similarly, we capture the base appearance of a target from a specific viewpoint, and then we can use text prompts to drive diffusion models [53, 21] for text-driven projection mapping. It is precisely because GS-ProCams efficiently models view-agnostic geometric and photometric mappings of ProCams that it can facilitate the application of ProCams in reality.
Figure 10: Additional visual comparisons of ProCams simulation on the Nepmap synthetic dataset. The first column displays an object from a novel viewpoint, the second column shows the object under a novel projection pattern, and the third and fourth columns present the results of two methods.
Figure 11: Additional visual comparisons of projector compensation on real-world benchmark dataset. Each row illustrates the results of a scene under a novel viewpoint that our proposed GS-ProCams has not previously seen. The other two view-specific methods necessitate repeated data capturing and training for these viewpoints. The first three columns respectively illustrate the surface appearance from the specified viewpoint, the uncompensated appearance, and the expected appearance. The last three columns display the compensated results from different models.
Appendix C More results
We present additional qualitative results of ProCams simulation and projector compensation with novel projections under novel viewpoints that GS-ProCams has not seen previously. Additional visual comparisons of ProCams simulation on the Nepmap synthetic dataset [14] are shown in Fig. 11. Additional visual comparisons of ProCams simulation on real-world benchmark dataset are given in Fig. 12. Additional comparisons of projector compensation on real-world benchmark dataset are provided in Fig. 11. Please note that the presented results are based on a novel viewpoint and a novel projection that GS-ProCams has not seen before within each trained scene.
Figure 12: Additional visual comparisons of ProCams simulation on real-world benchmark dataset. Each row illustrates the results of a target surface under a novel viewpoint that our proposed GS-ProCams has not previously seen, while DeProCams has retrained for it.